A BODY AT REST IS TO BE ACCELERATED BY A CONSTANT

FORCE SO THAT IT WILL HAVE A VELOCITY OF 800 CM/S

AFTER EXACTLY 8 S. FIND THE FORCE IF THE BODY WEIGHS 490

DYNES.


1
Expert's answer
2022-02-23T08:31:53-0500

Given:

"\Delta v=8\:\rm m/s"

"\Delta t=8\:\rm s"

"mg=490*10^{-5}\:\rm N"


The force

"F=m\frac{\Delta v}{\Delta t}\\
=(490*10^{-5}/9.8)*\frac{8}{8}=5*10^{-4}\:\rm N=50\: dynes"